### v1

# PRISM: Deriving the Transformer as a Signal-Denoising Operator via Maximum Coding Rate Reduction

Abstract
Deep learning models, particularly Transformers, are often criticized as ”black boxes” and lack interpretability. We propose Prism, a white-box attention-based architecture derived from the principles of Maximizing Coding Rate Reduction ($\text{MCR}^{2}$). By modeling the attention mechanism as a gradient ascent process on a distinct signal-noise manifold, we introduce two physical constraints: an overcomplete dictionary to expand the representational phase space, and an irrational frequency separation ($\pi$-RoPE) to enforce incoherence between signal and noise subspaces. We demonstrate that these geometric inductive biases can be viewed as a physical constraint and they are sufficient to induce unsupervised functional disentanglement alone. Using TinyStories as a controlled testbed for verifying spectral dynamics, we observe that Prism spontaneously specializes its attention heads into spectrally distinct regimes: low-frequency heads capturing long-range causal dependencies (signal) and high-frequency heads handling local syntactic constraints (noise). Our results suggest that interpretability and performance are not a trade-off, but can be unified through principled geometric construction.
Transformer, White-box Deep Learning
1 Introduction
The Transformer architecture (Vaswani et al., 2017) has established itself as the fundamental building block of modern artificial intelligence, exhibiting a remarkable power-law relationship between parameter scale and performance (Kaplan et al., 2020). However, despite their empirical success, these models operate largely as black boxes. One limitation is the architectural conflation between semantic reasoning (the ability to manipulate abstract concepts over long-range contexts) and syntactic memorization (the rote statistical completion of local patterns). While recent mechanistic interpretability efforts have identified specific circuits such as induction heads (Olsson et al., 2022), the standard attention mechanism lacks the physical inductive bias required to separate these distinct functional regimes by design. This may often result in brittle reasoning, hallucination driven by overfitting to noise and data inefficiency.
To enhance interpretability, we turn to first-principles approaches that view deep learning as data compression. The principle of Maximal Coding Rate Reduction ($MCR^{2}$) (Chan et al., 2022) shows that the fundamental goal of representation learning is to expand the volume of the dataset globally while compressing the volume of similar classes locally. Building on this theoretical foundation, the CRATE (Coding-RATE transformer) framework (Yu et al., 2024) establishes that the self-attention mechanism is mathematically isomorphic to a step of gradient ascent optimizing the rate reduction. This perspective transforms the attention layer mechanism from an empirical heuristic into a derived operator for signal denoising on low-dimensional manifolds, providing the necessary white-box substrate for our architectural interventions.
However, gaps between white-box neural networks and black-box neural networks still exist. One critical gap lies in data modeling; current white-box models typically assume clean data or Gaussian noise, whereas real-world data distributions are highly non-Gaussian. This may result in the coherence of optimization trajectory. If the model has limited capacity, the gradient flows of semantic signals (typically low-frequency global dependencies) and syntactic noise (high-frequency local statistics) mix together and hinder the learning. To mitigate this, we propose Prism, a physics-informed transformer architecture that introduces two specific geometric constraints to the standard Transformer block based on explicit additive noise. First, we impose an overcomplete dictionary (Candès et al., 2006) to expand the representational phase space, allowing signal and noise to separate in higher dimensions without competing for the same limited basis vectors. Second, we implement frequency separation via $\pi$-scaled Rotary Positional Embeddings ($\pi$-RoPE) (Su et al., 2024). By scaling the rotational bases of the overcomplete subspaces by $\pi$ and $1/\pi$ respectively, we hypothesize that this choice effectively mitigates resonance and keeps the reasoning and memorization dynamics on spectrally disjoint manifolds.
We empirically validate the Prism architecture under controlled constraints using a compact model with 50M parameters trained on the TinyStories benchmark (Eldan and Li, 2023). Despite the restricted scale, our experiments suggest that Prism reveals a phenomenon of unsupervised functional emergence and good training stability. We observe that the attention heads spontaneously organize into spectrally distinct regimes: heads driven by the low-frequency basis specialize in tracking long-range correlations (e.g. semantic signal), while those driven by the high-frequency basis tend to capture local correlations (e.g. syntactic structure). These results provide a proof-of-concept that disentangled reasoning capabilities can be induced via principled geometric construction rather than solely through massive parameter scaling.
2 Theoretical Framework
2.1 Rate Reduction as the Objective
We begin with the assumption that a useful signal $\mathbf{Z}$ can be understood as the combination of compressible useful $\mathbf{Z}_{0}$ and the noise $\mathbf{E}$, which yields:
$\begin{split}\mathbf{Z}=\mathbf{Z}_{0}+\mathbf{E},\end{split}$ (1)
and the learning objective is to extract $\mathbf{Z}_{0}$ from $\mathbf{Z}$. To achieve this, we follow the maximum coding rate reduction framework. Let $\mathbf{Z}\in\mathbb{R}^{N\times d}$ denote representation (e.g. input tokens), and $\mathbf{U}_{s},\mathbf{U}_{e}$ be the subspace with respect to the effective signal $\mathbf{Z}_{0}$ and noise $\mathbf{E}$ respectively. We assume the signal and noise reside in distinct subspaces, and these subspaces $\mathbf{U}_{s}$ and $\mathbf{U}_{n}$ are sufficiently incoherent (e.g., $\left\|\mathbf{U}_{s}^{\top}\mathbf{U}_{n}\right\|\approx 0$). We can approximate the noise $\mathbf{E}=\mathbf{U}_{n}^{\top}\mathbf{Z}$. Thus, the effective signal $Z_{0}$ is obtained by maximizing the coding rate difference $\Delta R(\mathbf{Z})$:
$\mathbf{Z}_{0}=\max_{\mathbf{Z}}\Delta R(\mathbf{Z})=R(\mathbf{Z}|\mathbf{U}_{s})-\lambda R(\mathbf{Z}|\mathbf{U}_{n}),$ (2)
where $R(\mathbf{Z})$ represents the coding rate of the signal projected onto local subspaces expanded by $\mathbf{U}_{s}$, and $R(\mathbf{Z}|\mathbf{U}_{n})$ represents the coding rate of the noise obtained by projecting the noisy signal onto the local subspaces $U_{n}$, $\lambda>0$ is the parameter controlling the separation between signal/noise decomposition. Specifically, under the Gaussian assumption, the coding rate of the matrix $\mathbf{Z}$ is defined as:
$R(\mathbf{Z})=\frac{1}{2}\log\det\left(\mathbf{I}+\frac{d}{N\epsilon^{2}}\mathbf{Z}\mathbf{Z}^{\top}\right),$ (3)
where $\epsilon$ is the distortion of encoding.
If there are many subspaces $\mathbf{U}=[\mathbf{U}_{1},\dots,\mathbf{U}_{K}]$, the coding rate of $\mathbf{Z}$ is the sum of the coding rates of all subspace projections, and yields
$R(\mathbf{Z}\mathbf{U}_{k})=\frac{1}{2}\log\det\left(\mathbf{I}+\beta(\mathbf{Z}\mathbf{U}_{k})^{\top}(\mathbf{Z}\mathbf{U}_{k})\right),$ (4)
where $\beta=\frac{p}{N\epsilon^{2}}$ and $p$ is the dimension of subspaces.
2.2 Prism as Gradient Ascent of Coding Rate Reduction
Figure 1: The PRISM White-Box Architecture. (Top) The input latent state $\mathbf{Z}^{\ell}$ is projected into an overcomplete feature space via dictionary $\mathbf{U}=[\mathbf{U}_{s},\mathbf{U}_{n}]$. The data flow in prism bifurcates into Signal Stream (Blue) and Noise Stream (Orange). $\pi$-RoPE: the Signal stream applies low-frequency rotary embeddings scaled by $\pi$ ($e^{i\pi\theta}$) to represent long-range semantic structures, while the Noise stream applies high-frequency embeddings ($e^{i\pi^{-1}\theta}$) to capture short-range syntactic artifacts. The layer output is computed via a differential operator, $\mathbf{Z}^{\ell+1}\leftarrow S-\lambda N$, where $\lambda$ is an annealing coefficient that dynamically suppresses the noise subspace, encouraging the model to learn long-term correlation (e.g. causal logic) during training.
To optimize the representations $\mathbf{Z}$ and remove the noise layer-by-layer, we apply a single step of gradient ascent on the objective $\Delta R(\mathbf{Z})$ with step size $\eta$. This yields the dynamic update rule similar to the Transformer block:
$\mathbf{Z}^{\ell+1}=\mathbf{Z}^{\ell}+\eta\frac{\partial\Delta R(\mathbf{Z}^{\ell})}{\partial\mathbf{Z}^{\ell}}.$ (5)
The Prism block is obtained by deriving the gradients for the log-determinant terms. Following (Yu et al., 2024), the derivative of the log-determinant term yields the softmax structure. We recover the structure of the one-head self-attention mechanism under the white-box constraint $\mathbf{Q}=\mathbf{K}=\mathbf{V}=\mathbf{U}$ and the output matrix is set directly into $\mathbf{W}_{O}=\mathbf{U}^{\top}$:
$\begin{split}\mathbf{Z}^{\ell+1}\approx&\mathbf{Z}^{\ell}+\eta\underbrace{\text{softmax}\left(\frac{(\mathbf{Z}^{\ell}\mathbf{U}_{s})(\mathbf{Z}^{\ell}\mathbf{U}_{s})^{\top}}{\sqrt{d}}\right)}_{\text{Attention Matrix}}(\mathbf{Z}^{\ell}\mathbf{U}_{s})\mathbf{U}^{\top}_{s}\\ &-\lambda\eta\underbrace{\text{softmax}\left(\frac{(\mathbf{Z}^{\ell}\mathbf{U}_{n})(\mathbf{Z}^{\ell}\mathbf{U}_{n})^{\top}}{\sqrt{d}}\right)}_{\text{Attention Matrix}}(\mathbf{Z}^{\ell}\mathbf{U}_{n})\mathbf{U}_{n}^{\top}.\end{split}$ (6)
Here, the term $\mathbf{Z}^{\ell}\mathbf{U}_{s}$ represents the projection onto the subspace $\mathbf{U}_{s}$, effectively removing the noise if the noise is incoherent with the useful signal.
An illustration of the Prism block is shown in Fig. 1, the input signal is first projected and separated and then modulated by $\pi$-RoPE to capture different features of different scale, and finally processed via differential de-noising to remove the noise.
For multi-head attention, we just introduce subspaces $\mathbf{U}_{s,1},\dots,\mathbf{U}_{s,K}]$ and $\mathbf{U}_{n,1},\dots,\mathbf{U}_{n,K}]$ for clean signal and noise respectively. Under the same procedure, we can obtain the multi-head attention directly:
$\begin{split}\mathbf{Z}^{\ell+1}&\approx\mathbf{Z}^{\ell}+\eta\sum_{k}\left\{\mathop{\text{Att}}(\mathbf{U}_{s,k},\mathbf{Z})(\mathbf{Z}^{\ell}\mathbf{U}_{s,k})\mathbf{U}^{\top}_{s,k}\right.\\ &~~~\left.-\lambda\mathop{\text{Att}}(\mathbf{U}_{n,k},\mathbf{Z})(\mathbf{Z}^{\ell}\mathbf{U}_{n,k})\mathbf{U}_{n,k}^{\top}\right\},\end{split}$ (7)
where $\mathop{\text{Att}}(\cdot,\cdot)$ is the attention matrix.
In practice, the underlying subspaces $\mathbf{U}_{s},\mathbf{U}_{n}$ are unknown and must be learned end-to-end via backpropagation. To eliminate the variance explosion, and considering the fact that the dictionary can be overcomplete, we suggest choosing $\eta=\frac{1}{\sqrt{RK}}$ where $R$ is the expansion ratio of the dictionary and input space  [^1] In standard Transformer the ratio is $1$, and we suggest setting $R=2$ to ensure same number of parameters between Prism and standard Transformer.. From this perspective, we can view weight optimization as finding the subspaces $\mathbf{U}_{s}$ and $\mathbf{U}_{n}$ maximizing the coding rate reduction and separate noise from noisy signal.
2.3 Mean-Field Dynamics and Spectral Decoupling
Considering a one-layer network with Prism block, as the number of heads $H\to\infty$, the summation becomes integral and the output of the Prism layer reduces to
$f(z)=z+\int\Phi(u,z)\rho_{S}(u)d(u)-\lambda\int\Phi(u,z)\rho_{N}(u)du$ (8)
where $\rho_{S}(u)$ and $\rho_{N}(u)$ are the probability density function of the weights $u$, and the kernel function $\Phi(x,z)$ is the kernel with respect to attention matrix. Although practical models use a finite number of heads, the mean-field formulation provides insights into the asymptotic behavior of the signal/noise separation dynamics.
Consider the mean square error with data $y$,
$\mathcal{L}(\rho_{S},\rho_{N})\approx\frac{1}{2}\mathbb{E}_{z}|y-(\text{Signal}-\lambda\cdot\text{Noise})|^{2},$
under the mean-field limit the Wasserstein gradient flow is given by:
[…]
$~~~\left.\left.-\lambda\int K_{SN}d\rho_{N}\right)\right)$
[…]
$~~~+\left.\left.\lambda^{2}\int K_{NN}d\rho_{N}\right)\right)$
where $K(u,v)=\int\Phi(u,z)\Phi(v,z)\rho_{z}dz$ is the kernel function characterizing the coherence or similarity between different heads and the $V_{\mathrm{data}}^{S}(u)$ and $V_{\mathrm{data}}^{N}(u)$ are terms with respect to training data, and $K_{SN},K_{NN},K_{SN}$ are the cross-interaction terms between signal and noise heads.
In sequential modeling, the time step $t$ is usually embedded by RoPE. Under RoPE, each head has its own frequency. Using the properties of RoPE $\langle R_{t}q,R_{t-\tau}k\rangle=q^{\top}R_{\tau}k$ where $R_{t}$ is the rotation matrix defined by RoPE and $\tau$ is the time step interval. Motivated by this, we thus decouple the pre-softmax logit for head $u$ by semantic part and position part:
$S_{u}(\tau)=C_{u}\cdot\cos(\omega_{u}\tau+\phi_{u}),$ (9)
where $C_{u}$ is some function depending on $u$ representing content. Due to the fact that the position and absolute content are usually decoupled, and each output is auto-regressive thus we approximate the kernel function by:
$K(u,v)\sim\mathbb{E}[C_{u}C_{v}]\cdot\mathbb{E}_{\tau}[\cos\omega_{u}\tau\cdot\cos\omega_{v}\tau],$ (10)
where $\mathbb{E}_{\tau}$ is the integral of different step $\tau$.
The second term can be written as:
$\cos\omega_{u}\tau\cdot\cos\omega_{v}\tau=\frac{1}{2}\cos((\omega_{u}+\omega_{v})\tau)+\cos((\omega_{u}-\omega_{v})\tau)$ (11)
The first term is oscillatory and vanishes as $\tau\xrightarrow{\infty}$ by Riemannian-Lebesgue lemma. Resonance occurs when the frequencies between signal and noise heads align, leading to coupling. To eliminate this, we want the orbit to be dense and uniformly distributed on the torus. Motivated by Weyl’s Equidistribution Theorem  [^2] https://en.wikipedia.org/wiki/Equidistribution_theorem, we suggest choosing the irrational frequency in RoPE. However, mere irrationality may be insufficient; we require not only non-resonance but also that the system is far from resonance. This requirement is analogous to the stability conditions in Hamiltonian mechanics, specifically the Kolmogorov-Arnold-Moser (KAM) theorem  [^3] https://en.wikipedia.org/wiki/Kolmogorov-Arnold-Moser_theorem. In the series expansion of the interaction kernel $K_{SN}$, denominators of the form $\frac{1}{n\omega_{S}-m\omega_{N}}$ appear (where $n,m$ are integers). If this denominator is too small, the system becomes unstable. So we finally suggest using irrational numbers satisfying Diophantine conditions in RoPE  [^4] https://en.wikipedia.org/wiki/Diophantine_approximation The numbers are very hard to approximate by rational numbers.. This leads to our $\pi$-RoPE [^5] One may choose $\phi=\frac{1+\sqrt{5}}{2}$, but the number may be too small to separate the long-range and short-range signal.
$\theta_{S}=\pi\theta_{\mathrm{base}}\quad\theta_{N}=\frac{1}{\pi}\theta_{\mathrm{base}},$ (12)
where $\theta_{\mathrm{base}}$ is the base frequency [^6] Usually denoted as $b$. in RoPE that is usually chosen to be $10000$ in GPT-2. More theoretical benefit can be shown, if we study the Wasserstein gradient flow of Prism with respect to $\rho_{S}$ and $\rho_{N}$, their evolution equation has an interaction kernel term which can be eliminated by $\pi$-RoPE. Thus, the subspace between signal and noise are nearly decoupled and the whole optimization becomes a good optimization problem. The signal heads find the long-range and low-frequency signal and the noise heads fit the high-frequency, short-range solution.
Finally we adopt $\pi$-RoPE as an ansatz to design our rotary embeddings and give a conjecture below:
Conjecture : (Non-Resonance Condition in Prism). We posit that selecting a rotational frequency ratio $\omega_{S}/\omega_{N}\notin\mathbb{Q}$ (an irrational number, specifically Diophantine Numbers like $\pi$) minimizes the integrated cross-correlation $K_{\text{SN}}$ between the Signal and Noise subspaces over long sequences.
3 Experiments
To validate the theoretical hypothesis of signal-noise decomposition, we conducted experiments on the TinyStories dataset (Eldan and Li, 2023) , which is a benchmark designed to test reasoning and coherence in small-scale language models. We deliberately choose a compact setup to conduct controlled intervention analysis, isolating the geometric mechanism from the interference of massive parameter redundancy.
Model Configuration: We implemented a ”Prism-mini” variant and a GPT-2 baseline. We train a 8-layer Prism with overcomplete attention ($R=2$, resulting in $16$ physical heads per layer). The heads are split into pairs of signal head and noise head
RoPE Configuration: We applied $\pi$-RoPE for frequency separation. Signal heads utilized a base frequency $\theta_{S}=10000\pi\approx 31415$, while Noise heads utilized $\theta_{N}=10000/\pi\approx 3183$.
Training Details: We implement the Prism via nanoGPT  [^7] https://github.com/karpathy/nanoGPT/tree/master and then train the Prism-mini from scratch on a single NVIDIA RTX 4090 GPU for $20000$ steps with a batch size of $32$ and context length of $512$. We utilized the AdamW optimizer with cosine learning rate decay. The learning rate is 6e-4, and the warmup step is $1000$. We use the cosine schedule from $0.01$ to $0.1$ for hyperparameter $\lambda$ to stabilize training.
Figure 2: Training Dynamics Emergent Functional Specialization (a) Training and validation loss curves on the TinyStories dataset. (b) Evolution of the L2 Gradient Norm. (c)-(d) Comparative visualization of attention maps in Layer 6, Head 2. (c) Signal Head: Driven by low-frequency $\pi$-RoPE, it captures sparse, long-range semantic dependencies (e.g., attending to specific entities). (d) Noise Head: Driven by high-frequency $\pi$-RoPE, it captures local syntactic artifacts and background noise.
The training dynamics and the visualization of attention matrix is shown in Fig. 2. The Fig. 2 (a) indicates that Prism (50M) exhibits rapid convergence, achieving a validation loss of $\approx 1.55$ and surpassing the GPT-2 (22M) baseline (red dashed line  [^8] https://huggingface.co/jonasknobloch/gpt2_u020_tiny-stories_1024_dpos), demonstrating high parameter efficiency even better than Gemma 3 (dashed lines in blue, purple, and green  [^9] https://github.com/di37/gemma3-270M-tinystories-pytorch). Despite the potential instability of the differential operator ($S-\lambda N$), Fig. 2 (b) shows that the training is stable, and the gradient norm stabilizes at a low magnitude ($\approx 0.5$), indicating a smooth optimization landscape facilitated by the white-box constraints. Specific Visualization in Fig. 2 (c) and Fig. 2 (d) indicates that the signal heads of Prism automatically focus on the long-range semantic link (such as the causal link between ”door” and ”key”), whereas the noise heads focus more on local syntax (e.g. dependency between ”for” and ”door”).
Figure 3: Overcomplete Signal Subspace Visualization: (a) Subject-centric: Maintains a persistent vertical attention band on the word “frog”, tracking the narrative protagonist. (b) Object-centric (Causal): When predicting the target token “door” (last row), this head explicitly attends to the distal causal antecedent “key”, bridging the logical gap between the tool and the target. (c) Structural: Attends to the determiner “a”, likely tracking the introduction of new entities. (d) Action-centric: Anchors the sequence to the primary predicate “found” and “looked”, tracing the action.
To visualize the learned signal, we visualize the attention patterns of four distinct heads of the sixth layer from the signal head group ($\pi$-scaled RoPE) for the input sequence “The frog found a shiny key… looked for a door”. As shown in Fig. 3, despite sharing the same low-frequency inductive bias, the overcomplete basis ($R=2$) enables the emergence of diverse semantic roles without supervision and reveals the richness within the signal subspaces. Thus, the overcomplete parameterization allows the model to allocate dedicated heads for specific semantic roles such as subject and action, effectively revealing a sparse dictionary learning process.
4 Conclusion, Discussion and Future Works
In this work, we derived the Prism layer from the first principle of maximizing coding rate reduction, demonstrating that semantic reasoning and syntactic processing are spectrally separable regimes within the Transformer architecture. By imposing an overcomplete dictionary and enforcing spectral separation via $\pi$-RoPE, Prism achieves unsupervised functional disentanglement, evidenced by the emergence of distinct, specialized functional heads.
Limitations of this study include the restricted parameter scale (50M). However, the Prism architecture retains high compatibility with modern computational primitives such as FlashAttention (Dao et al., 2022) and sparse scaling mechanisms like Mixture of Experts (MoE) (Shazeer et al., 2017). We are currently extending the framework to larger-scale datasets and deeper parameterizations to verify the persistence of these geometric properties at scale.
Theoretically, future work will focus on establishing a rigorous derivation of the mean-field dynamics under the overcomplete setting and focus on establishing bounds for the non-resonance condition for the transcendental $\pi$-RoPE basis. Furthermore, we aim to extend the white-box paradigm to the Feed-Forward Networks (FFN) via sparse coding unrolling. We believe that the FFN layer can be understood and replaced by a finite-step approximation of sparse coding optimization, effectively refining the signal by projecting it onto a union of low-dimensional subspaces.
Implications for Large Scale Models: We hypothesize that the ”Attention Sink” phenomena in current Large Language Models fundamentally stem from the spectral entanglement between global semantic signals and local syntactic noise. By enforcing irrational frequency separation via $\pi$-RoPE, Prism geometrically encourages that long-range causal dependencies remain orthogonal to high-frequency artifacts, preventing noise accumulation over extended contexts. Consequently, this suggests that structural geometric constraints, rather than mere data scaling, offer a first-principles path to resolving stability issues in 100100B+ parameter models.
Finally, we believe that de-noising and compression are the two fundamental imperatives in neural architecture design. Our results suggest that Prism, which is a white-box Transformer constructed from geometric principles, is not merely a modification of the standard Transformer, but a mathematically grounded alternative that achieves competitive performance while offering intrinsic interpretability. Our results indicate that interpretability and performance are not necessarily trade-offs, but can be unified through principled geometric construction, where functional specialization emerges naturally from rigorous physical constraints.
Acknowledgments
This work was supported by the Postdoctoral Fellowship Program of CPSF (Grant No. GZC20232943).